OpenStudy (anonymous):

5x-2y=2 graph or just value of y

OpenStudy (muzzack):

could be both

OpenStudy (muzzack):

we could first isolate y in just value of y

OpenStudy (muzzack):

and in graphing same thing

OpenStudy (muzzack):

so ttry it

OpenStudy (muzzack):

i cant really help if you wont try it

OpenStudy (muzzack):

i cant give out answers so plz

OpenStudy (texaschic101):

5x - 2y = 2 to find x intercept, sub in 0 for y 5x - 2(0) = 2 5x = 2 x = 2/5.....x intercept is (2/5,0) to find y intercept, sub in 0 for x 5(0) - 2y = 2 -2y = 2 y = -1....y intercept is (0,-1) to find the slope, put the equation in y = mx + b form 5x - 2y = 2 -2y = -5x + 2 y = 5/2x - 1 slope is 5/2 plot your intercepts...(2/5,0) and (0,-1) start at y intercept (0,-1) and since slope is 5/2, go up 5 spaces, then to the right 2 spaces, then up 5 and to the right 2, and you should cross the x axis at (2/5,0)